OpenStudy (anonymous):

A clock loses 2 min every 6hrs. At 3:00pm, the clock is set to the correct time and allowed to run with out interference. A. what integer would describe the time loss after exactly 3 days? B. what would the clock read at 3:00pm three days later?

OpenStudy (anonymous):

12*2 = loss

OpenStudy (anonymous):

since in three days there are24*3 hours =72 hours

OpenStudy (anonymous):

for A?

OpenStudy (anonymous):

and there are 12 intervals of 6 hours so loss= 2*12

OpenStudy (anonymous):

yes that's for A

OpenStudy (anonymous):

so the answer to a is 12*2

OpenStudy (anonymous):

yes 12*2

OpenStudy (anonymous):

after three days clock will show 3.00 p.m- 24 mins= 2 p.m and 38 miutes

OpenStudy (anonymous):

how did u get that one?

OpenStudy (anonymous):

since loss= 24 minutes therefore after three days at 3 p.m subtract 24 minutes from 3 p.m it was expected to be 3 p.m but it is not ...there is loss of 24 minutes

OpenStudy (anonymous):

ohh ok i get it now

OpenStudy (anonymous):

cool :)